\begin{references}
\bibitem{urs-err}
	U. A. Wiedemann, Phys. Rev. C {\bf 61} (2000) 029902E
\bibitem{urs-mpbe}
	U. A. Wiedemann, Phys. Rev. C {\bf 57} (1998) 3324
\bibitem{cstjz}
	T. Cs\"org\H o and J. Zim\'anyi, hep-ph/9705433,
	Phys. Rev. Lett. {\bf 80}, 916 (1998)
\bibitem{balazs}
	J. Zim\'anyi, T. Cs\"org\H o, B. Luk\'acs, M. Rhoades-Brown
	and N. L. Balazs, preprint KFKI - 1991 -28/A, p. 91.
\bibitem{jzcst}
	J. Zim\'anyi and T. Cs\"org\H o,
	hep-ph/9705432, Heavy Ion Phys. {\bf 9}, 241 (1999)
\bibitem{plcoh}
	T. Cs\"org\H o,
	quant-ph/9903050, Heavy Ion Phys. {\bf 9}, 161 (1999)
\bibitem{cs-nato}
	T. Cs\"org\H o,
	hep-ph/0001233, to appear in {\it ``Particle Production Spanning
	MeV and TeV Energies"}, Proc. {\bf NATO Advanced Study Institute}
	(Kluwer, 2000, ed. W. Kittel, P. Mulders and O. Scholten) p. 203.
\end{references}

\documentstyle[prl,aps]{revtex}

\begin{document}
\draft
\title{Comment on the Erratum: Multiparticle Bose-Einstein Correlations\\
	$\left[\right.$ Phys. Rev. C 61, 029902(E) (2000) of
	Phys. Rev. C 57, 3324 (1998) $\left. \right]$
	}

\author{T. Cs\"org\H o$^{1,2}$
	\thanks{E-mails: csorgo@sunserv.kfki.hu,
	csorgo@nt1.phys.columbia.edu}{\ }
}
\address{
{$^1$Department of Physics, Columbia University,
 538 W 120-th Street, New York, NY 10027 }\\
{$^2$MTA KFKI RMKI, H-1525 Budapest 114. POB. 49, Hungary}
 }
\date{\today}
\maketitle
\pacs{25.75.Gz,24.10.Cn,52.60.+h,99.10.+g}

Although the Erratum~\cite{urs-err} correctly
states that the solution of model (4.1) depends only on two
effective parameters, in contrast to earlier statements of
ref.~\cite{urs-mpbe},
the Erratum presents the formulas for
$R_{{\rm eff}}$ and $\Delta_{{\rm eff}}$ incorrectly,
as can be seen even from a dimensional analysis.
The correct formulas for model (4.1)
were given in eq. (16) of ref. ~\cite{cstjz},
using a notation $R_{{\rm eff}} = R_e$ and
$\Delta_{{\rm eff}} = \sigma_T$,
as introduced in ref.~\cite{balazs}.
The analytic solution of model (4.1) was summarized in ref.~\cite{cstjz},
detailed investigations of the model properties
were described in refs.~\cite{jzcst,plcoh,cs-nato}.

Supported by the US - Hungarian Joint Fund MAKA 652/1998,
by the OTKA grants T025435 and T029158, and by the
US Department of Energy Contracts No.
DE - FG02 -93ER40764, DE-FG-02-92-ER40699 and
DE - AC02 -76CH00016.

\end{document}